# DocReward: A Document Reward Model for Structuring and Stylizing Junpeng Liu1 Yuzhong Zhao211footnotemark: 1Bowen Cao1Jiayu Ding3Yilin Jia4Tengchao Lv5 Yupan Huang5Shaohan Huang5Nan Yang5Li Dong5Lei Cui5Tao Ge5Xun Wang5 Huitian Jiao5Sun Mao5FNU Kartik5Si-Qing Chen5Wai Lam1Furu Wei5 1CUHK 2UCAS

### v1

@@ 4.2 Baselines and Evaluation Settings
Table 4: Extrinsic evaluation results. DocReward shows utility for professional document generation.
4.4 Improving Document Generation with DocReward
To demonstrate the effect of our DocReward as a reward model in document generation, we conduct a extrinsic evaluation. A document agent generates documents given the same text content and then a reward model identifies the best one from the documents according to their scores. We compare three reward models: random, GPT-5, and DocReward. Human annotators rank the selected documents from each reward models according to their structure and style. Finally, we calculate the win/lose/tie rates for each reward model against the others. As presented in Table 4, the random baseline performs poorly, winning only of comparisons and losing . GPT-5 achieves more balanced results with a win rate of . By contrast, DocReward substantially outperforms both baselines, achieving a win rate of and losing only of the time. These results indicate that DocReward’s reward signal better captures the structural and stylistic qualities that humans value. The evaluation demonstrates that plugging DocReward into a standard document agent improves the final, human-preferred output without changing the underlying agent itself. The evaluation details are presented in Appendix A.5.
4.5 Case Study
We present a case study on documents with identical textual content but different structures and styles in Figure 5. In case (a), the allocation of whitespace is ineffective, with insufficient space between Last Name and excessive space between First Name, leading to an imbalanced layout. Key fields such as Faculty/Department, Country, and Country Code are not vertically aligned, causing a cluttered and disorganized layout. This poor alignment and inconsistent spacing result in a low score of 1.21 from DocReward. Case (b) adopts a table-like arrangement, but the level-1 heading The teaching staff member is too small and does not stand out from the body text, diminishing its impact. Additionally, the lack of borders around input fields makes it hard to locate items easily, resulting in a moderate score of 2.11. Case (c) provides a clear and well-structured layout, with headings appropriately larger than the body text and better readability, earning the highest rating of 5.34. These results show that DocReward effectively captures document professionalism in structure and style. Additional cases are provided in Appendix A.7.
(a) score: 1.21
(b) score: 2.11
(c) score: 5.34
Figure 5: DocReward’s assessment of structural and stylistic professionalism.
@@ 4.6 Visualization of Attention Map
To understand DocReward’s internal decision-making process, we conduct probing experiments analyzing its attention maps within the language model part. The attention maps are computed over image patches. As shown in Figure 6, the attention maps reveal that the model relies more on structural and formatting cues than on semantic content when evaluating document professionalism. In Figure 6, attention is focused on headings and numbering, indicating sensitivity to structure clarity and logical flow. The model also allocates considerable attention to page headers (i.e., “CS-66”) and footers at bottom right corner (i.e., “DEC. 2006”), suggesting that the inclusion of page headers and footers is an important signal of professional structure. In Figure 6, the model attends strongly to bullet points, suggesting that formatting consistency and emphasis markers are key professionalism signals. In Figure 6, attention is dispersed across table grids, highlighting the importance of text alignment and readability in structured tabular layouts. Moreover, the attention maps show notable focus on the four page corners, suggesting that DocReward implicitly checks for uniform margins and balanced whitespace, which are strong indicators of professional layout design.
@@ 5 Related Work
Additionally, photo aesthetics are modeled using layout-aware CNNs such as A-Lamp (Li et al., 2018), and similar techniques extend to video (Liu & Yu, 2023).
These studies show that aesthetic principles can guide AI development and that human preferences are reliable supervisory signals, but they focus on images or UI interfaces rather than multi-page documents, where professionalism depends on both structure and style.
Document AI: Structure and Generation.
@@
This demonstrates that preference data offers a scalable and effective way to align generative models with nuanced expectations.
@@ 6 Conclusion
In this paper, we introduced DocReward, a document reward model designed to assess structural and stylistic professionalism. Our key contributions include the construction of a multi-domain dataset DocPair of 117K paired documents, each with high- and low-professionalism counterparts. We train DocReward using the Bradley-Terry loss. Rigorous evaluation on a human-annotated test set demonstrated DocReward’s superior performance, outperforming GPT-4o and GPT-5 by 30.6, 19.4 percentage points, respectively in human preference accuracy.
Moreover, a human preference evaluation demonstrates its utility to guide generation agents toward producing human-preferred documents.
@@
Shuai Bai, Keqin Chen, Xuejing Liu, Jialin Wang, Wenbin Ge, Sibo Song, Kai Dang, Peng Wang, Shijie Wang, Jun Tang, et al.
@@
arXiv preprint arXiv:2502.13923, 2025.
Bradley & Terry (1952)
Ralph Allan Bradley and Milton E Terry.
@@
Biometrika, 39(3/4):324–345, 1952.
@@
https://github.com/comfyanonymous/ComfyUI, 2025.
@@
Simon R Davies, Richard Macfarlane, and William J Buchanan.
@@
Forensic Science International: Digital Investigation, 40:301330, 2022.
@@
Shubhang Shekhar Dvivedi, Vyshnav Vijay, Sai Leela Rahul Pujari, Shoumik Lodh, and Dhruv Kumar.
@@
In Proceedings of the 1st ACM international conference on AI-powered software, pp. 65–73, 2024.
@@
Yuzheng Fu, Yulin Zhang, Yuchen Li, Ruoyu Zheng, and Qiming Liu.
@@
IEEE Transactions on Multimedia, 2024.
@@
Simson Garfinkel, Paul Farrell, Vassil Roussev, and George Dinolt.
@@
Sirui Hong, Mingchen Zhuge, Jonathan Chen, Xiawu Zheng, Yuheng Cheng, Ceyao Zhang, Jinlin Wang, Zili Wang, Steven Ka Shing Yau, Zijuan Lin, et al.
@@
International Conference on Learning Representations, ICLR, 2024.
@@
Aaron Hurst, Adam Lerer, Adam P Goucher, Adam Perelman, Aditya Ramesh, Aidan Clark, AJ Ostrow, Akila Welihinda, Alan Hayes, Alec Radford, et al.
@@
arXiv preprint arXiv:2410.21276, 2024.
@@
Zhouqiang Jiang, Bowen Wang, Junhao Chen, and Yuta Nakashima.
@@
arXiv preprint arXiv:2410.10471, 2024.
@@
Chenglong Li, Jiansheng Wang, Song Yan, Chen Song, and Lianyi He.
@@
In Proceedings of the 26th ACM international conference on Multimedia, pp. 177–185, 2018.
@@
Yuxin Li, Qinglin Zhang, Chen Fang, Jin Xiao, and Wen Xu.
@@
In The Eleventh International Conference on Learning Representations (ICLR), 2023.
Liang et al. (2025)
Xinbin Liang, Jinyu Xiang, Zhaoyang Yu, Jiayi Zhang, Sirui Hong, Sheng Fan, and Xiao Tang.
Openmanus: An open-source framework for building general ai agents, 2025.
URL https://doi.org/10.5281/zenodo.15186407.
@@
Jiawei Lin, Jiaqi Guo, Shizhao Sun, Zijiang Yang, Jian-Guang Lou, and Dongmei Zhang.
@@
Advances in Neural Information Processing Systems, 36:43852–43879, 2023.
Liu & Yu (2023)
Chang Liu and Han Yu.
@@
Damiano Marsili, Rohun Agrawal, Yisong Yue, and Georgia Gkioxari.
@@
In Proceedings of the Computer Vision and Pattern Recognition Conference (CVPR), pp. 19446–19455, June 2025.
@@
Long Ouyang, Jeffrey Wu, Xu Jiang, Diogo Almeida, Carroll Wainwright, Pamela Mishkin, Chong Zhang, Sandhini Agarwal, Katarina Slama, Alex Ray, et al.
@@
Advances in neural information processing systems, 35:27730–27744, 2022.
@@
Sida Peng, Eirini Kalliamvakou, Peter Cihon, and Mert Demirer.
@@
arXiv preprint arXiv:2302.06590, 2023.
@@
Rafael Rafailov, Archit Sharma, Eric Mitchell, Christopher D Manning, Stefano Ermon, and Chelsea Finn.
@@
Advances in neural information processing systems, 36:53728–53741, 2023.
@@
Long Ouyang Stiennon, Jeffrey Wu, Daniel M Ziegler, Ryan Lowe, Jeremy Voss, Alec Radford, Dario Amodei, and Paul F Christiano.
@@
Advances in Neural Information Processing Systems, 33:11345–11358, 2020.
@@
Nishant Subramani, Alexandre Matton, Malcolm Greaves, and Adrian Lam.
@@
arXiv preprint arXiv:2011.13534, 2020.
@@
Zecheng Tang, Chenfei Wu, Juntao Li, and Nan Duan.
@@
arXiv preprint arXiv:2309.09506, 2023.
@@
Jiaxu Tian, Xuehui Yu, Yaoxing Wang, Pan Wang, Guangqian Guo, and Shan Gao.
@@
arXiv preprint arXiv:2507.05568, 2025.
@@
Jian Xie, Kai Zhang, Jiangjie Chen, Tinghui Zhu, Renze Lou, Yuandong Tian, Yanghua Xiao, and Yu Su.
@@
arXiv preprint arXiv:2402.01622, 2024.
@@
Yiheng Xu, Minghao Li, Lei Cui, Shaohan Huang, Furu Wei, and Ming Zhou.
@@
In Proceedings of the 26th ACM SIGKDD international conference on knowledge discovery & data mining, pp. 1192–1200, 2020.
Yan et al. (2025)
Yibo Yan, Shen Wang, Jiahao Huo, Philip S. Yu, Xuming Hu, and Qingsong Wen.
MathAgent: Leveraging a mixture-of-math-agent framework for real-world multimodal mathematical error detection.
In Georg Rehm and Yunyao Li (eds.), Proceedings of the 63rd Annual Meeting of the Association for Computational Linguistics (Volume 6: Industry Track), pp. 69–82, Vienna, Austria, July 2025. Association for Computational Linguistics.
ISBN 979-8-89176-288-6.
doi: 10.18653/v1/2025.acl-industry.7.
URL https://aclanthology.org/2025.acl-industry.7/.
@@
ACM Transactions on Information Systems (TOIS), 37(4):1–27, 2019.
@@
Xueru Zhang, Chen Peng, Xin Liao, Shugong Hou, Yang Fan, Li Liang, Yaohui Zhao, and Chunhui Zhang.
@@
In The Twelfth International Conference on Learning Representations (ICLR), 2024.
@@
Yaowei Zheng, Richong Zhang, Junhao Zhang, Yanhan Ye, Zheyan Luo, Zhangchi Feng, and Yongqiang Ma.
@@
arXiv preprint arXiv:2403.13372, 2024.
@@
Ziwei Zheng, Michael Yang, Jack Hong, Chenxiao Zhao, Guohai Xu, Le Yang, Chao Shen, and Xing Yu.
Deepeyes: Incentivizing” thinking with images” via reinforcement learning.
arXiv preprint arXiv:2505.14362, 2025.
[…]
@@ A.2 Model Implementation Details
Our document reward model is built upon the Qwen2.5-VL multimodal architecture, with the maximum input pixels set to 300,000. It is configured with a maximum context length of 16,000 tokens to ensure comprehensive understanding. Training utilizes the AdamW optimizer with a learning rate of 1e-6 and a batch size of 256 over 3 epochs. All training was conducted on 8 NVIDIA A100 GPUs. The training code is based on LLaMA-Factory (Zheng et al., 2024).

### v2

@@ 4.3 Improving Document Generation with DocReward
Table 3: Best-of-N evaluation results. DocReward shows utility for professional document generation.
Reward Types
Success
ROUGE-L
DocReward
Rank
Qwen2.5-Coder
30.0
20.61
0.0663
4.58
+ rule
98.0
97.94
0.1785
4.06
+ DocReward
100.0
97.95
0.3046
2.84
GPT-4o
52.0
48.73
0.2682
3.18
+ rule
66.0
62.15
0.3189
2.70
+ DocReward
78.0
74.33
0.4486
2.02
Table 4: Results of the reinforcement learning experiments. “DocReward” denotes the sigmoid-normalized DocReward score. “Rank” denotes the average ranking assigned by human annotators.
Best-of-N.
A document agent generates candidates from the same textual content, and a reward model selects the best one based on its score. We compare three reward models including “Random”, GPT-5 and DocReward with human annotators ranking the selected outputs by structure and style. As shown in Table 4, DocReward consistently outperforms the baselines, indicating that its reward better aligns with human preferences. This evaluation shows that integrating DocReward into a document agent improves output quality without modifying the agent itself. Details are provided in Appendix A.7.
Reinforcement Learning.
We aim to enhance document generation agentic workflows that take plain text as input and generate professional documents. We consider two kinds of rewards: 1) that penalizes documents that either result from invalid python-docx code or differ from the input text after execution. Specifically, if the code executes successfully, then ; else will be zero. 2) that penalizes documents with poor structure and style. Overall, the reward assigned to each generated document is defined as:
where is a hyperparameter to balance the rewards, is the operation to regularize the value range of DocReward to , and represents whether is larger than a threshold.
For open-source models, we adopt GRPO (Shao et al., 2024) as the reinforcement learning algorithm, while employing training-free GRPO (Cai et al., 2025) for closed-source models. After RL, human annotators are asked to rank the documents produced by six model variants. The evaluation criterion is the professionalism of the document’s structure and styling. The experimental results are shown in Table 4. Rule-based rewards substantially improve the success rate of document generation for both Qwen2.5-Coder and GPT-4o. Incorporating DocReward as reward further enhances the performance of both open- and closed-source models, leading to higher success rates, improved ROUGE-L scores, and better average rankings. These results demonstrate that DocReward serves as an effective reward model for professional structure and style. Figure 5 presents visualization of documents generated by different models.
Figure 5: Visualization of documents generated by various models. “Execution error” denotes instances where the generated code encountered runtime failures, preventing the final document from being rendered.
Table 6: Cross-lingual robustness evaluation. denotes the performance gap between English and non-English scenarios. DocReward demonstrates strong cross-lingual robustness.
Out-of-Domain Evaluation. Table 6 reports the out-of-domain results of different models. Firstly, DocReward-7B (77.5) remains superior to all baseline models, including the closed-source model GPT-5 (68.4). This trend is consistent with the in-domain results. Secondly, the performance of DocReward-7B decreases by merely 4.8 percentage points when transitioning from in-domain to out-of-domain evaluations. Such a small performance gap indicates that DocReward generalizes effectively to unseen domains.
Cross-Lingual Robustness: To evaluate robustness across languages, we conduct experiments in French, Spanish, and Russian, with results shown in Table 6. Firstly, in non-English scenarios, DocReward-7B model achieved a remarkable score of 77.9, substantially outperforming all baseline models. Secondly, all models, including the baselines, exhibited performance degradation in non-English settings. The performance drop of DocReward () is even smaller than that of closed-source models GPT-4o () and GPT-5 (), indicating that DocReward demonstrates strong cross-lingual robustness.
@@ 4.5 Visualization of Attention Map
To analyze DocReward’s decision-making process, we conduct probing experiments on attention maps. As shown in Figure 6, the model attends more to structural and formatting cues than semantic content when assessing document professionalism. Specifically, attention focuses on headings and numbering (Figure 6), page headers and footers (e.g., “CS-66”, “DEC. 2006”), bullet points (Figure 6), and table grids (Figure 6), reflecting sensitivity to structure clarity, formatting consistency, alignment, and readability. Additionally, attention to page corners suggests that uniform margins and balanced whitespace are important indicators of professional layout design.
@@ 5 Related Work
Additionally, photo aesthetics are modeled using layout-aware CNNs such as A-Lamp (Li et al., 2018), and similar techniques extend to video (Liu and Yu, 2023).
These studies focus on images or UI interfaces rather than multi-page documents, where professionalism depends on both structure and style.
Document AI.
@@ 6 Conclusion
In this paper, we introduced DocReward, a document reward model designed to assess structural and stylistic professionalism.
To enable professionalism assessment without being influenced by textual quality, we propose a textual-quality-agnostic framework for document reward modeling. Then, a multi-domain preference dataset, DocPair, consisting of 117K paired documents, is constructed to enable the training of DocReward using the Bradley-Terry loss.
Comprehensive experiments demonstrate that DocReward not only outperforms GPT-5 in assessing structure and style, but also serves as an effective reward model in reinforcement learning for agentic workflows.
7 Limitations
A primary limitation is that DocReward is designed as a scalar reward model, providing only a numerical score to represent the overall structural and stylistic quality of a document. It currently cannot generate natural language feedback explaining why a document is perceived as unprofessional. Therefore, extending DocReward from a scalar model to an interpretable reward model that can produce both quantitative scores and qualitative diagnostic rationales remains a key focus of our future work.
@@ 7 Limitations
Shuai Bai, Keqin Chen, Xuejing Liu, Jialin Wang, Wenbin Ge, Sibo Song, Kai Dang, Peng Wang, Shijie Wang, Jun Tang, and 1 others. 2025.
@@
arXiv preprint arXiv:2502.13923.
Bradley and Terry (1952)
Ralph Allan Bradley and Milton E Terry. 1952.
@@
Biometrika, 39(3/4):324–345.
Cai et al. (2025)
Yuzheng Cai, Siqi Cai, Yuchen Shi, Zihan Xu, Lichao Chen, Yulei Qin, Xiaoyu Tan, Gang Li, Zongyi Li, Haojia Lin, and 1 others. 2025.
Training-free group relative policy optimization.
arXiv preprint arXiv:2510.08191.
@@
https://github.com/comfyanonymous/ComfyUI.
@@
Simon R Davies, Richard Macfarlane, and William J Buchanan. 2022.
@@
Forensic Science International: Digital Investigation, 40:301330.
@@
Shubhang Shekhar Dvivedi, Vyshnav Vijay, Sai Leela Rahul Pujari, Shoumik Lodh, and Dhruv Kumar. 2024.
@@
In Proceedings of the 1st ACM international conference on AI-powered software, pages 65–73.
@@
Yuzheng Fu, Yulin Zhang, Yuchen Li, Ruoyu Zheng, and Qiming Liu. 2024.
@@
IEEE Transactions on Multimedia.
@@
Simson Garfinkel, Paul Farrell, Vassil Roussev, and George Dinolt. 2009.
@@
Sirui Hong, Mingchen Zhuge, Jonathan Chen, Xiawu Zheng, Yuheng Cheng, Ceyao Zhang, Jinlin Wang, Zili Wang, Steven Ka Shing Yau, Zijuan Lin, and 1 others. 2024.
@@
International Conference on Learning Representations, ICLR.
Hui et al. (2024)
Binyuan Hui, Jian Yang, Zeyu Cui, Jiaxi Yang, Dayiheng Liu, Lei Zhang, Tianyu Liu, Jiajun Zhang, Bowen Yu, Keming Lu, Kai Dang, Yang Fan, Yichang Zhang, An Yang, Rui Men, Fei Huang, Bo Zheng, Yibo Miao, Shanghaoran Quan, and 5 others. 2024.
Qwen2.5-coder technical report.
Preprint, arXiv:2409.12186.
@@
Aaron Hurst, Adam Lerer, Adam P Goucher, Adam Perelman, Aditya Ramesh, Aidan Clark, AJ Ostrow, Akila Welihinda, Alan Hayes, Alec Radford, and 1 others. 2024.
@@
arXiv preprint arXiv:2410.21276.
@@
Zhouqiang Jiang, Bowen Wang, Junhao Chen, and Yuta Nakashima. 2024.
@@
arXiv preprint arXiv:2410.10471.
@@
Chenglong Li, Jiansheng Wang, Song Yan, Chen Song, and Lianyi He. 2018.
@@
In Proceedings of the 26th ACM international conference on Multimedia, pages 177–185.
@@
Yuxin Li, Qinglin Zhang, Chen Fang, Jin Xiao, and Wen Xu. 2023.
@@
In The Eleventh International Conference on Learning Representations (ICLR).
@@
Jiawei Lin, Jiaqi Guo, Shizhao Sun, Zijiang Yang, Jian-Guang Lou, and Dongmei Zhang. 2023.
@@
Advances in Neural Information Processing Systems, 36:43852–43879.
Liu and Yu (2023)
Chang Liu and Han Yu. 2023.
@@
Damiano Marsili, Rohun Agrawal, Yisong Yue, and Georgia Gkioxari. 2025.
@@
In Proceedings of the Computer Vision and Pattern Recognition Conference (CVPR), pages 19446–19455.
@@
Long Ouyang, Jeffrey Wu, Xu Jiang, Diogo Almeida, Carroll Wainwright, Pamela Mishkin, Chong Zhang, Sandhini Agarwal, Katarina Slama, Alex Ray, and 1 others. 2022.
@@
Advances in neural information processing systems, 35:27730–27744.
@@
Sida Peng, Eirini Kalliamvakou, Peter Cihon, and Mert Demirer. 2023.
@@
arXiv preprint arXiv:2302.06590.
@@
Rafael Rafailov, Archit Sharma, Eric Mitchell, Christopher D Manning, Stefano Ermon, and Chelsea Finn. 2023.
@@
Advances in neural information processing systems, 36:53728–53741.
Shao et al. (2024)
Zhihong Shao, Peiyi Wang, Qihao Zhu, Runxin Xu, Junxiao Song, Xiao Bi, Haowei Zhang, Mingchuan Zhang, YK Li, Yang Wu, and 1 others. 2024.
Deepseekmath: Pushing the limits of mathematical reasoning in open language models.
arXiv preprint arXiv:2402.03300.
@@
Long Ouyang Stiennon, Jeffrey Wu, Daniel M Ziegler, Ryan Lowe, Jeremy Voss, Alec Radford, Dario Amodei, and Paul F Christiano. 2020.
@@
Advances in Neural Information Processing Systems, 33:11345–11358.
@@
Nishant Subramani, Alexandre Matton, Malcolm Greaves, and Adrian Lam. 2020.
@@
arXiv preprint arXiv:2011.13534.
@@
Zecheng Tang, Chenfei Wu, Juntao Li, and Nan Duan. 2023.
@@
arXiv preprint arXiv:2309.09506.
@@
Jiaxu Tian, Xuehui Yu, Yaoxing Wang, Pan Wang, Guangqian Guo, and Shan Gao. 2025.
@@
arXiv preprint arXiv:2507.05568.
@@
Jian Xie, Kai Zhang, Jiangjie Chen, Tinghui Zhu, Renze Lou, Yuandong Tian, Yanghua Xiao, and Yu Su. 2024.
@@
arXiv preprint arXiv:2402.01622.
@@
Yiheng Xu, Minghao Li, Lei Cui, Shaohan Huang, Furu Wei, and Ming Zhou. 2020.
@@
In Proceedings of the 26th ACM SIGKDD international conference on knowledge discovery & data mining, pages 1192–1200.
@@
ACM Transactions on Information Systems (TOIS), 37(4):1–27.
@@
Xueru Zhang, Chen Peng, Xin Liao, Shugong Hou, Yang Fan, Li Liang, Yaohui Zhao, and Chunhui Zhang. 2024.
@@
In The Twelfth International Conference on Learning Representations (ICLR).
@@
Yaowei Zheng, Richong Zhang, Junhao Zhang, Yanhan Ye, Zheyan Luo, Zhangchi Feng, and Yongqiang Ma. 2024.
@@
arXiv preprint arXiv:2403.13372.
@@
Ziwei Zheng, Michael Yang, Jack Hong, Chenxiao Zhao, Guohai Xu, Le Yang, Chao Shen, and Xing Yu. 2025.
Deepeyes: Incentivizing" thinking with images" via reinforcement learning.
arXiv preprint arXiv:2505.14362.
@@ A.2 Model Implementation Details
Our document reward model is built upon the Qwen2.5-VL multimodal architecture, and a regression head is added to predict scalar scores. The maximum input pixels are set to 300,000. It is configured with a maximum context length of 16,000 tokens. Training utilizes the AdamW optimizer with a learning rate of 1e-6 and a batch size of 256 over 3 epochs. All training was conducted on 8 NVIDIA A100 GPUs. The training code is based on LLaMA-Factory (Zheng et al., 2024).